# FlowBind: Efficient Any-to-Any Generation with Bidirectional Flows

### v1

@@ 1 Introduction
Bridging the gap from specialist to generalist flow models introduces fundamental hurdles, primarily due to requirements of multi-modal data and computational cost. Frameworks that rely on a central anchor modality, typically text (tang2023codi), requires each modality to be paired with text during training so that all modalities can be aligned through the shared text representation. This design is restrictive, as it prevents the model from learning the rich, direct correlations that exist beyond language. Conversely, methods that model the full joint conditioning of all modalities (li2025omniflow) can achieve expressive generation performance but at a steep cost: they require some fully-paired data for stable training, which is scarce, and their computational complexity often scales quadratically with the number of modalities. These data and compute issue render them impractical for real-world scenarios with a large and diverse set of modalities.
@@ 3 FlowBind
where $v^{i}$ denotes the modality-specific velocity field. Note that multi-modal flows are factorized per modality given the shared latent (Eq. 4), and the shared latent implicitly aligns these flows across modality (Eq. 5). During training, the shared latent is instantiated as $z^{*}=H_{\phi}(\mathbf{z}^{S})$ through an auxiliary encoder $H_{\phi}$, whose marginal approximates $\pi_{\text{shared}}$ and is optimized jointly with the per-modality drift networks $v_{\theta^{i}}$ (Figure 1(a)). At inference, FlowBind relies only on the learned drift networks: owing to the invertibility of the direct flows, both inferring the shared latent from input modalities and generating outputs from the latent are achieved by a single drift network per modality (Figure 1(b)). Details of the training and inference procedures are provided in Section 3.1.
@@ 4 Related Work
Another line of work utilizes discrete diffusion models, often by adapting them to operate on discrete token spaces (rojas2025diffuse_every; shi2025muddit). These methods, which typically focus on text-image generation tasks, leverage the high-quality synthesis capabilities of diffusion for multi-modal scenario. For instance, UniDisc (swerdlow2025unidisc) highlights the controllability of this approach by framing various conditional generation tasks, such as inpainting.
@@
Recent flow-based models (liu2025crossflow; he2025flowtok) have explored learning direct, data-to-data invertible mappings between two modalities, predominantly focusing on text-image pairs. This approach represents a fundamental departure from traditional generative flows that typically learn bridging from fixed prior distributions (e.g., standard Gaussian) to target data distributions through conditional generation mechanisms. To facilitate these direct transformations, existing methodologies designate latent distribution of one modality (i.e., source distribution) as a learnable embedding. This is achieved by introducing an encoder for the source modality and constructing additional loss terms that align the source and target modalities, such as contrastive learning objectives.
@@ 5.2 Results on One-to-One Generation
Category Model T →I $\text{I}\to\text{T}$ $\text{T}\to\text{A}$ $\text{A}\to\text{T}$ $\text{I}\to\text{A}$ $\text{A}\to\text{I}$
FID ↓ $\text{CIDEr}\uparrow$ $\text{FAD}\downarrow$ $\text{CIDEr}\uparrow$ $\text{FAD}\downarrow$ $\text{FID}\downarrow$
Specialists SD3-Medium $25.40$
FLUX.1 $22.06$
LLaVA-NeXT -- 109.3
TangoFlux -- 1.41
AudioX -- 3.09
Qwen2-Audio -- 4.64
Seeing & Hearing -- 5.31
Sound2Vision -- 42.55
@@
Category Model T →I $\text{I}\to\text{T}$ $\text{T}\to\text{A}$ $\text{A}\to\text{T}$ $\text{I}\to\text{A}$ $\text{A}\to\text{I}$
CLIP ↑ $\text{CLIP}\uparrow$ $\text{CLAP}\uparrow$ $\text{CLAP}\uparrow$ $\text{AIS}\uparrow$ $\text{AIS}\uparrow$
Specialists SD3-Medium $31.60$
FLUX.1 $31.06$
LLaVA-NeXT -- 32.14
TangoFlux -- 42.71
AudioX -- 29.29
Qwen2-Audio -- 17.09
Seeing & Hearing -- 75.11
Sound2Vision -- 62.39
@@
In terms of data efficiency, FlowBind is achieves any-to-any generation with much smaller training dataset, using 0.15 % of CoDi and 1.79 % of OmniFlow. We conjecture the training can be done with much smaller dataset because we choose to model flow between high-level representations. By doing so, the cross-modal generation capability is decomposed into inter-modal alignment and intra-modal generation in FlowBind. Our drift network is only required to capture inter-modal correspondence, as per-modality frozen encoder-decoders take charge in capturing intra-modal generative capability. This would enable FlowBind to quickly capture cross-modal alignment with smaller datasets.
@@ 5.3 Results on Many-to-Many Generation
Table 4: Many-to-one generation alignment performances.
(I+A) $\rightarrow$ T (T+A) $\rightarrow$ I (T+I) $\rightarrow$ A
[…]
CoDi 24.04 20.66 25.17 57.52 4.85 61.28
OmniFlow 26.38 36.07 24.06 54.90 7.68 59.32
[…]
Table 5: One-to-many generation alignment performances.

### v2

@@ 1 Introduction
Bridging the gap from specialist to generalist flow models introduces fundamental hurdles, primarily due to requirements of multi-modal data and computational cost. Frameworks that rely on a central anchor modality, typically text (tang2023codi), require each modality to be paired with text during training so that all modalities can be aligned through the shared text representation. This design is restrictive, as it prevents the model from learning the rich, direct correlations that exist beyond language. Conversely, methods that model the full joint conditioning of all modalities (li2025omniflow) can achieve expressive generation performance but at a steep cost: they require some fully-paired data for stable training, which is scarce, and their computational complexity often scales quadratically with the number of modalities. These data and compute issues render them impractical for real-world scenarios with a large and diverse set of modalities.
@@ 3 FlowBind
where $v^{i}$ denotes the modality-specific velocity field. Note that multi-modal flows are factorized per modality given the shared latent (Eq. 4), and the shared latent implicitly aligns these flows across modalities (Eq. 5). During training, the shared latent is instantiated as $z^{*}=H_{\phi}(\mathbf{z}^{S})$ through an auxiliary encoder $H_{\phi}$, whose marginal approximates $\pi_{\text{shared}}$ and is optimized jointly with the per-modality drift networks $v_{\theta^{i}}$ (Figure 1(a)). At inference, FlowBind relies only on the learned drift networks: owing to the invertibility of the direct flows, both inferring the shared latent from input modalities and generating outputs from the latent are achieved by a single drift network per modality (Figure 1(b)). Details of the training and inference procedures are provided in Section 3.1.
@@ 4 Related Work
Another line of work utilizes discrete diffusion models, often by adapting them to operate on discrete token spaces (rojas2025diffuse_every; shi2025muddit). These methods, which typically focus on text-image generation tasks, leverage the high-quality synthesis capabilities of diffusion for multi-modal scenarios. For instance, UniDisc (swerdlow2025unidisc) highlights the controllability of this approach by framing various conditional generation tasks, such as inpainting.
@@
Recent flow-based models have explored learning direct, data-to-data invertible mappings between two data distributions (li2023bbdm; wang2024semflow), predominantly focusing on text-image pairs (liu2025crossflow; he2025flowtok). This approach represents a fundamental departure from traditional generative flows that typically learn bridging from fixed prior distributions (e.g., standard Gaussian) to target data distributions through conditional generation mechanisms. To facilitate these direct transformations, existing methodologies designate latent distributions of one modality (i.e., source distribution) as a learnable embedding. This is achieved by introducing an encoder for the source modality and constructing additional loss terms that align the source and target modalities, such as contrastive learning objectives.
@@ 5.2 Results on One-to-One Generation
Category  Model  $\mathrm{T}\to\mathrm{I}$ $\mathrm{I}\to\mathrm{T}$  $\mathrm{T}\to\mathrm{A}$ $\mathrm{A}\to\mathrm{T}$  $\mathrm{I}\to\mathrm{A}$ $\mathrm{A}\to\mathrm{I}$
    $\mathrm{FID}\downarrow$ $\mathrm{CIDEr}\uparrow$  $\mathrm{FAD}\downarrow$ $\mathrm{CIDEr}\uparrow$  $\mathrm{FAD}\downarrow$ $\mathrm{FID}\downarrow$
Specialists  SD3-Medium    25.40 –    –
  FLUX.1    22.06 –    –
  LLaVA-NeXT    109.3    –
  TangoFlux    –    1.41 –
  AudioX    –    3.09 –
  Qwen2-Audio    –    4.64
  Seeing & Hearing    –    –    5.31
  Sound2Vision    –    –    42.55
@@
Category  Model  $\mathrm{T}\to\mathrm{I}$ $\mathrm{I}\to\mathrm{T}$  $\mathrm{T}\to\mathrm{A}$ $\mathrm{A}\to\mathrm{T}$  $\mathrm{I}\to\mathrm{A}$ $\mathrm{A}\to\mathrm{I}$
    $\mathrm{CLIP}\uparrow$ $\mathrm{CLIP}\uparrow$  $\mathrm{CLAP}\uparrow$ $\mathrm{CLAP}\uparrow$  $\mathrm{AIS}\uparrow$ $\mathrm{AIS}\uparrow$
Specialists  SD3-Medium    31.60 –    –
  FLUX.1    31.06 –    –
  LLaVA-NeXT    32.14    –
  TangoFlux    –    42.71 –
  AudioX    –    29.29 –
  Qwen2-Audio    –    17.09
  Seeing & Hearing    –    –    75.11
  Sound2Vision    –    –    62.39
@@
In terms of data efficiency, FlowBind achieves any-to-any generation with a much smaller training dataset, using 0.15 % of CoDi and 1.79 % of OmniFlow. We conjecture the training can be done with much smaller dataset because we choose to model flow between high-level representations. By doing so, the cross-modal generation capability is decomposed into inter-modal alignment and intra-modal generation in FlowBind. Our drift network is only required to capture inter-modal correspondence, as per-modality frozen encoder-decoders take charge in capturing intra-modal generative capability. This would enable FlowBind to quickly capture cross-modal alignment with smaller datasets.